Circular No. 8531

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 JQ_{5} (CATALINA) An apparently asteroidal object found by the Catalina Survey, and

designated 2005 JQ_{5} (cf. MPEC 2005-J29; discovery observation given below), has been found to show a bright asymmetric coma on seven 40-s co-added R-band CCD exposures taken by C. Snodgrass, S. C. Lowry, and A. Fitzsimmons with the 2.0-m Faulkes Telescope-North on May 17.5 UT. Earlier reports mentioned hints at “softness” of the object’s images (S. Larson, discovery frames with the 0.68-m Schmidt telescope) and a 1′-diameter very-low-surface-brightness coma (B. A. Skiff, LONEOS 0.59-m Schmidt telescope, May 11.3).

2005 UT  α_{2000}  δ_{2000}  Mag.
May 6.27858  15^{h}14^{m}09.^{s}92  −3^{o}10′44′′.3  16.7

Additional astrometry, elliptical orbital elements (T = 2005 July 28.1 TT, q = 0.826 AU, ω = 222^{o}.7, Ω = 95^{o}.8, i = 5^{o}.7 (equinox 2000.0), e = 0.690, P = 4.35 yr), and an ephemeris appear on MPEC 2005-K14.

2QZ J142701.6–012310 E. Rykoff, Physics Department, University of Michigan, on behalf of

the ROTSE collaboration, reports that ROTSE-III sky-patrol images taken on Jan. 23 revealed an outburst of 2QZ J142701.6–012310 (initially classified as a possible blazar in the 2dF quasar survey by Véron-Cetty and Véron 2003, A.Ap. 412, 399) at mag R = 15.1; the variable was not detected in ROTSE-III images taken on Jan. 21 (R > 18.0). The object is visible as a very blue pointlike object in the Sloan Digital Sky Survey at R = 20.4. Rykoff adds that a spectrum taken on Jan. 25 with the Hobby-Eberly Telescope at McDonald Observatory shows no obvious absorption or emission features.

P. A. Woudt and B. Warner, University of Cape Town (UCT), report that 2QZ J142701.6–012310 appears to be a new AM CVn-type cataclysmic variable. High-speed photometry during quiescence (V ∼ 20.3), obtained with the South African Astronomical Observatory 1.9-m telescope (+ UCT CCD camera) on May 15 and 16 has revealed a photometric modulation at 2194 ± 4 s, typical of an AM CVn-type variable. Spectroscopy obtained in quiescence (Croom et al. 2004, MNRAS 349, 1397) and during its January outburst (Rykoff et al., above) shows that the spectrum appears rich in helium, consistent with it being a double-degenerate helium-transferring binary (only the twelfth-known such object).

2005 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green